Knowledge Check 02

Pastel Papers Inc. produces wrapping paper. The standard direct materials quantity is 0.20 pound per unit and the standard rate is $0.30 per pound. During June, the company actually purchased and used 1,800 pounds of direct materials to produce 7,500 units, at an average cost of $0.40 per pound. Calculate the direct materials price variance.

  • $180 favorable

  • $750 favorable

  • $750 unfavorable

  • $180 unfavorable

Pastel Papers Inc. produces wrapping paper. The standard direct materials quantity is 0.20 pound per unit and the standard rate is $0.30 per pound. During June, the company actually purchased and used 1,800 pounds of direct materials to produce 7,500 units, at an average cost of $0.40 per pound. Calculate the direct materials quantity variance.

  • $90 unfavorable

  • $60 unfavorable

  • $90 favorable

  • $60 favorable

Calculate the direct materials spending variance when the direct materials price variance is $5,500 favorable and the direct materials quantity variance is $660 unfavorable.

  • $4,840 unfavorable

  • $6,160 unfavorable

  • $4,840 favorable

  • $6,160 favorable
Answer #1

1) Material price variance = (0.30-0.40)*1800 = 180 U

So answer is d) $180 U

2) Material quantity variance = (7500*.2-1800)*0.30 = 90 U

So answer is a) $90 U

3) Material spending variance = 5500-660 = 4840 F

So answer is c) $4840 Favorable